OpenStudy (anonymous):

lim x approaches 1 |(x-1)|/(x-1)

OpenStudy (aravindg):

What did you get when you tried to evaluate this?

OpenStudy (anonymous):

0? i plugged in 1 but i dont think thats correct..

OpenStudy (aravindg):

See first convert modulus into function at different intervals |x|=x when x>0 |x|=-x when x<0 Write this for |x-1|

OpenStudy (anonymous):

don't be hoodwinked by this

OpenStudy (anonymous):

\[\frac{|x-1|}{x-1}\] is only ever one of two numbers if \(x-1>0\) then it is \(1\) and if \(x-1<0\) then it is \(-1\) try a few numbers and you will see it instantly

OpenStudy (anonymous):

so i just plug in numbers greater and less than 1?

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

what i meant was, if you plug in any number larger than 1 for x, you will get 1

OpenStudy (anonymous):

\[\frac{|3-1|}{3-1}=\frac{2}{2}=1\] for example

OpenStudy (anonymous):

so for any number larger than 1, this expression is just the number 1

OpenStudy (anonymous):

and for any number less than one, this expression is just the number \(-1\) for example, if \(x=0\) you get \[\frac{|0-1|}{0-1}=\frac{1}{-1}=-1\]

OpenStudy (anonymous):

if x - 1 > and < aren't equal, does that mean the limit doesn't exist?

OpenStudy (anonymous):

in other words \[\frac{|x-1|}{x-1} = \left\{\begin{array}{rcc}- 1& \text{if} & x <1 \\ 1& \text{if} & x>1 \end{array} \right. \]

OpenStudy (anonymous):

and yes, since the limit from the right is 1 and from the left is -1, that means the limit does not exist, like you said

OpenStudy (anonymous):

ok thank you i understand

OpenStudy (anonymous):

yw